If someone makes $8 per hour and time and a half for any hours over 40, how much do they make if they work 45 hours one week?
Answer by checkley71(8403)

40*8+(1.5*8)5=320+12*5=320+60=$380 FOR A 45 HOUR WEEK WITH TIME & A HALF FOR THE LAST 5 HOURS.